Fractions: If the cost of the science fair is $300, how much should each school pay based on the number of students? MM 100 students WM 60 students TM 40 students
Answer by edjones(8007)

300/200=3/2=1.50 per student
100*1.50=$150.
60*1.50=$90.
40*1.50=$60.
.
Ed